How much energy must the shock absorbers of a 1150-kg car dissipate in order to damp a bounce that initially has a velocity of 0.4 m/s at the equilibrium position? Assume the car returns to its original vertical position.

J
Answer #1

Energy absorbed by the shock absorbers are

E=(1/2)mV2 =(1/2)*1150*0.42

E=92 J